OpenStudy (anonymous):

Solve the quadratic equation: 3x2 − 2x − 8 = 0 A.4/3,2 B.-4/3,3 C. -2/3,4 D. -4/3,2

OpenStudy (e.cociuba):

Do u mean: 3x^2 − 2x − 8 = 0 ??

OpenStudy (anonymous):

ya

OpenStudy (e.cociuba):

(3x + 4) (x - 2) = 0 x = -4/3 x = 2 So, your answers gunna be D

OpenStudy (e.cociuba):

Do u want me to explain?

OpenStudy (anonymous):

yes plz

OpenStudy (e.cociuba):

So basically u want to factor the problem out to get to this: (3x + 4) (x - 2) = 0 Do u know how to factor?

OpenStudy (anonymous):

no i do not know how to factor thing really good

OpenStudy (e.cociuba):

In order to factor 3x^2 − 2x − 8 = 0, first multiply the leading coefficient 3 and the last term -8 to get -24.

OpenStudy (e.cociuba):

So here comes the factoring part: What two numbers multiply to -24 and add to -2?

OpenStudy (e.cociuba):

So the answers gunna be 4 and -6.

OpenStudy (e.cociuba):

Do u have any questions so far?

OpenStudy (anonymous):

no

OpenStudy (e.cociuba):

Ok so..the original quadratic 3x^2 − 2x − 8, breaks down to this (just replace -2x with the two numbers that multiply to -24 and add to -2, which are: 4 and -6):